OpenStudy (anonymous):

Solve the equation. 1/a=6/18

OpenStudy (anonymous):

\[\frac{ 1 }{ a }=\frac{ 6 }{ 18 }\]

OpenStudy (whpalmer4):

Cross-multiply: 1*18 = a*6 Now can you solve for \(a\)?

OpenStudy (anonymous):

\[Is this the answer \frac{ 1 }{ 3 }=\frac{ 6 }{ 18 }\]

OpenStudy (whpalmer4):

Yes, because \[\frac{1}{3}=\frac{1}{3}*\frac{6}{6} = \frac{1*6}{3*6}=\frac{6}{18}\]